# Rolling with the Punches: Resilient Contrastive Pre-training under Non-Stationary Drift

### v2

@@ Abstract
The evolution of large-scale contrastive pre-training propelled by top-tier datasets has reached a transition point in the scaling law. Consequently, sustaining and enhancing a model’s pre-training capabilities in drift environments have surfaced as a notable challenge. In this paper, we initially uncover that contrastive pre-training methods are significantly impacted by concept drift wherein distributions change unpredictably, resulting in notable biases in the feature space of the pre-trained model. Empowered by causal inference, we construct a structural causal graph to analyze the impact of concept drift to contrastive pre-training systemically, and propose the causal interventional contrastive objective. Upon achieving this, we devise a resilient contrastive pre-training approach to accommodate the data stream of concept drift, with simple and scalable implementation. Extensive experiments on various downstream tasks demonstrate our resilient contrastive pre-training effectively mitigates the bias stemming from the concept drift data stream. Codes are available at https://anonymous.4open.science/r/ResilientCL/.
Machine Learning, ICML
@@ 1 Introduction
Contrastive learning has proven to be highly effective in pre-training large-scale models, especially in large vision models exemplified by frameworks like SimCLR (chenSimpleFrameworkContrastive2020; chenBigSelfSupervisedModels2020), MoCo series (heMomentumContrastUnsupervised2020; chenEmpiricalStudyTraining2021), DINO series (caronEmergingPropertiesSelfSupervised2021; zhouDINOWMWorldModels2024). However, with the ongoing scaling of large models, data hunger for contrastive learning is raising more attention in the community towards pre-training effectively from drift data. It could be caused by long-tailed data, noise, and domain shift, where concept drift (luLearningConceptDrift2019; yang2024adaptingmultimodallargelanguage) is utilized to uniformly summarize this phenomenon of unpredictable distribution changes in the pre-training through contrastive learning. Hence, a pertinent question emerges: beyond the existing contrastive learning methods, can contrastive paradigm learn from drift pre-training? In this work, we aim to bridge this gap by providing a systematic analysis of the above question. Our findings highlight critical vulnerabilities of the current contrastive pre-training paradigm in adapting to these challenges, underscoring the need for novel strategies to enhance their robustness in drift data streams. More related works are provided in Appendix A.
Current contrastive pre-training methods predominantly adhere to the paradigm of comparing two distinct views of the same object, typically derived from different encoders. Especially in terms of large vision models, student-teacher structures are used to extract different features from two views of images for contrasting agreement, like DINO (caronEmergingPropertiesSelfSupervised2021), MoCo v3 (chenEmpiricalStudyTraining2021). The student network is propelled by the contrastive loss, whereas the teacher network is momentum updated through an exponential moving average (EMA) of the student parameters. In the context of drift pre-training, this contrastive manner will be subtly affected by the concept drift of data streams and be accumulated through momentum, leading to amplified bias. We utilize long-tailed drift as a representative case to demonstrate the impact of concept drift on contrastive pre-training. Illustrated in Fig. 1, we visualize the feature space to investigate how long-tailed drift impacts the contrastive pre-training of MoCo v3 (chenEmpiricalStudyTraining2021). ImageNet (russakovskyImagenetLargeScale2015) and ImageNet-LT (liuLargeScaleLongTailedRecognition2019) are separately leveraged as balanced and long-tailed sources for contrastive pre-training, where the test data for generating feature space is balanced. Categories are differentiated by varying colors ranging from dark to light, indicating a spectrum from categories with few training samples to those with abundant samples. It is evident from Fig. 1(b), under imbalanced drift pre-training, the tail categories are confined to a small portion of the feature space, whereas the majority is dominated by the head categories with ample samples. The model is steered away from the center of balanced feature space, due to the accumulated bias in momentum update.
@@
(b) Long-Tailed Drift Pre-training
@@
Under the above assumption, we formulate the causal relationships (pearl2014interpretation) among samples $X$, predictions $Y$, the concept drift $D$ within data streams and the bias $B$ in the pre-trained teacher network with momentum update in a causal graph as shown in Fig.2(a). We find the concept drift $D$ within data streams is a confounder of $X$ and $Y$, where the backdoor (pearl1995causal) path $X\leftarrow D\rightarrow B\rightarrow Y$ leads to the false correlation, and the mediation (pearl2022direct) path $X\rightarrow B\rightarrow Y$ confounds the direct contribution of $X\rightarrow Y$. More analyses and details are given in Section 2.1.
Therefore, we propose a novel contrastive learning-based pre-training method with causal intervention (pearl2016causal) to mitigate the bias caused by concept drift in data streams.
In summary, our paper mainly makes the following contributions:
@@
    We are pioneers in uncovering the causal effects of concept drift on contrastive pre-training, especially in the accumulated bias through momentum update. It will facilitate future research in conducting a more comprehensive examination of concept drift on contrastive pre-training and leveraging larger datasets for pre-training large models.
@@
    By constructing the structural causal graph, we leverage causal inference theory to systemically analyze the concept drift within the biased datastream under the contrastive objective. With the aforementioned discussion, interventional contrastive pre-training is proposed to disentangle the drift effect during the pre-training. It offers a straightforward and scalable implementation for more abundant datasets and larger models.
@@
    Extensive experiments evaluate the performance of our contrastive method under drift pre-training. In the context of downstream tasks of long-tailed classification, OOD detection and domain shift, ours demonstrates superior performance with results of fine-tuning and linear probing. Crucially, we assess the inter-category distances to evaluate the efficacy of pre-training, where our model outperforms others, demonstrating our model effectively addresses drift in contrastive pre-training, facilitating large-scale pre-training.
@@ 2 Methodology
2.1 Causal Views of Drift Pre-training
Concept drift is a statistical phenomenon wherein the distributional characteristics of a target domain undergo arbitrary changes over the course of training (luLearningConceptDrift2019). Regarding the drift pre-training, we formula these unpredictable distribution changes with data stream $S_{0,t}=\{s_{0},...,s_{t}\}$, where $s_{i}=(x_{i},z_{i})$ represents a sample with feature vectors $x_{i}$ and contrastive label $z_{i}$, and $t$ denotes the timestamp in the training data stream, corresponding to the iteration step of the pre-training process. In iteration $t$, $S_{0,t}$ follows a certain distribution $F_{0,t}(x,z)$, thus the drift pre-training can be formalized as:
$\exists t:P_{t}(x,z)\neq P_{t+1}(x,z)$ (1)
where the joint probability $P_{t}(x,z)$ can be decomposed as $P_{t}(x,z)=P_{t}(x)\times P_{t}(z|x)$.
Regarding the concept drift in contrastive learning, the momentum update of the teacher network can be expressed as the following formula:
$\theta^{m}_{t}=\lambda\theta^{m}_{t-1}+(1-\lambda)\theta^{g}_{t-1}$ (2)
where $\theta^{m}$ signifies the parameters of the teacher network, $\theta^{g}$ denote those of the student network, and $\lambda\in[0,1)$ represents the momentum coefficient. Momentum $\lambda\theta^{m}_{t-1}$ induces a slower evolution of the teacher network $\theta^{m}$ relative to the student network $\theta^{s}$. While this approach contributes to the stabilization of training, it can lead to the accumulation of biases that are not readily rectified, especially the large momentum advised by MoCo v3 (e.g. $\lambda=0.999$).Hence, in order to tailor contrastive learning for drift pre-training, it is crucial to adjust the model to synchronize with the evolving data distribution for maximum agreement, formally defined as
$\max_{g^{(t)},g^{(t+1)},...,g^{(t+\tau)}}\sum_{i=t}^{t+\tau}L(g^{(i)}(\tilde{x}^{i}),m^{(i)}(\hat{x}^{i}))$ (3)
where $\hat{x}$ and $\tilde{x}$ symbolize different augmented samples obtained from $x$, $g_{\theta}^{(t)}$ denotes the student model trained by the data stream $S_{t-k,t-1}$ from the drift adaption window with the size of $k$, and $m_{\theta}^{(t)}$ represents the momentum model. And the model is driven by the target metric $L$ continuously to adapt the drift in a given drift adaptation window of $[t,t+\tau]$ time period.
(a) Structural Causal model
@@ 2.1 Causal Views of Drift Pre-training
For analyzing contrastive learning under the drift pre-training, we construct a structural causal graph (pearl1995causal; pearl2016causal) to formula the causal relationship among sample features ($X$), prediction (Y), the contrastive teacher network ($T$) and the drift in the pre-training ($D$) as illustrated in Fig.2, where $A\rightarrow B$ denotes that $A$ is the causer of $B$. The causal graph of $\{X,Y,D,B\}$ presents the following causal connections:
@@
In the constructed causal graph, nodes $D$ and $B$ are identified as the confounder and the mediator (pearl2022direct), respectively. The confounder $D$ influences both the independent variable $X$ and the outcome variable $Y$ from the backdoor path $X\leftarrow D\rightarrow B\rightarrow Y$, thereby distorting the estimated effect of the sample features on the predictions and resulting in the spurious correlation. For example, in the case of long-tailed drift pre-training, tail samples may be incorrectly associated with head categories through this backdoor path, a misalignment attributable to this spurious correlation. Besides, mediator $B$ conveys the drift effect of sample features $X$ on predictions $Y$ via the pathway $X\rightarrow B\rightarrow Y$, weakening the direct impact of $X$ on $Y$.
@@ 2.2 Causal Interventional Contrastive Objective
Drawing from the aforementioned causal relationships, it can be argued that current contrastive pertaining approaches involving the InfoNCE loss (oordRepresentationLearningContrastive2019) essentially leverages the likelihood $P(Y|X)$ to drive the whole network, which is vulnerable by the concept drift confounder $D$ and leads to the emergence of spurious correlations. $P(Y|X)$ can be formulated as:
$P_{t}(Y|X)=\sum_{i}^{|D|}P_{t}(Y|X,B=h(X,d_{i}))P_{t}(d_{i}|X)$ (4)
where $h(X,d_{i})$ abstractly denotes the sample bias $B$ is a combination from $X$ and $d_{i}$.

### v3

# Resilient Contrastive Pre-training under Non-Stationary Drift

@@ Abstract
The remarkable success of large-scale contrastive pre-training has been largely driven by by vast yet static datasets. However, as the scaling paradigm evolves, this paradigm encounters a fundamental challenge when applied to dynamic data streams characterized by concept drift—unpredictable changes in the underlying data distribution. This paper aims to advance robust pre-training under such non-stationary environments. We begin by revealing that conventional contrastive pre-training methods are highly susceptible to concept drift, resulting in significant substantial bias and instability within the learned feature representations. To systematically analyze these effects, we develop a structural causal model that elucidates how drift acts as a confounder, distorting the learned representations. Based on these causal insights, we propose Resilient Contrastive Pre-training (RCP), a novel method that incorporates causal intervention. RCP formulates a causally-informed objective to mitigate drift-induced biases through targeted interventions. The method is designed for simple and scalable implementation and exhibits notable adaptability, promoting robust and autonomous pre-training on non-stationary data. Comprehensive experiments across various downstream tasks consistently demonstrate that RCP effectively alleviates the detrimental impact of concept drift, yielding more resilient and generalizable representations. Codes are available at https://anonymous.4open.science/r/ResilientCL/.
@@ 1 Introduction
Contrastive learning has emerged as a dominant paradigm for large-scale visual pre-training, demonstrating impressive performance across a wide range of vision tasks [heMomentumContrastUnsupervised2020, caronEmergingPropertiesSelfSupervised2021, zhouDINOWMWorldModels2024]. While most existing studies evaluate contrastive methods on static, fully-scraped internet corpora, the pre-training landscape is shifting toward dynamic and continuously expanding data regimes[yang2024adaptingmultimodallargelanguage, yang2025learning]. Modern large-scale systems no longer rely on a one-shot frozen dataset. Instead, they accumulate data over time from heterogeneous sources such as user-generated content in federated/edge-learning pipelines[gao2022feddc, chenimportance, panchal2023flash], medical imaging networks [jiang2022harmofl, chen2024general, yang2022local, yang2024segmentation, yang2021variational], and autonomous-driving fleets [liang2022effective, zheng2023autofed]. In these real-world settings, data arrive sequentially, exhibit strong non-IID characteristics, and naturally undergo concept drift, including evolving class frequencies, shifting domains, seasonal changes, and sensor-induced feature heterogeneity [luLearningConceptDrift2019, yang2024adaptingmultimodallargelanguage, yang2025walking]. Recent studies in federated learning [chenimportance, jiang2022harmofl, panchal2023flash] and autonomous driving [zheng2023autofed, liang2022effective] consistently highlight that drifting data distributions degrade representation quality and destabilize pre-trained backbones. These developments raise a critical but under-explored question:
Can contrastive pre-training reliably learn from non-stationary and drifting data streams?
Through theoretical analysis and empirical evidence, we show that current contrastive objectives are highly sensitive to temporal distribution shift, leading to biased momentum teachers, collapsed similarity structures, and unstable representation dynamics. This exposes an urgent need for principled approaches that make contrastive pre-training robust to drift in realistic, continuously evolving environments. A review of related work is provided in the Appendix.
Current contrastive pre-training methods primarily rely on maximizing the agreement between two augmented views of the same input instance, typically processed through different encoders or network states [chenSimpleFrameworkContrastive2020, grillBootstrapYourOwn2020, heMomentumContrastUnsupervised2020]. In the context of large vision models, this paradigm is often instantiated through student–teacher architectures, as exemplified by frameworks such as DINO  [caronEmergingPropertiesSelfSupervised2021] and MoCo v3 [chenEmpiricalStudyTraining2021]. In these setups, the student network is optimized via a contrastive objective, while the teacher network is updated using an exponential moving average (EMA) of the student parameters to stabilize training. However, this widely adopted momentum-based update mechanism becomes particularly vulnerable under concept drift. As the underlying data distribution evolves, the teacher network—updated only gradually through EMA—tends to retain outdated knowledge from previous distributions. This stale information accumulates over time, propagating outdated representations and amplifying the misalignment between the learned feature space and the current data characteristics, ultimately degrading the quality of the pre-trained representations.
@@
(b) Tailed Drift Pre-training
@@
To concretely illustrate this phenomenon, we examine the effect of long-tailed drift—a prevalent form of concept drift where class frequencies evolve over time—on the contrastive pre-training process. We visualize the feature space learned by MoCo v3 [chenEmpiricalStudyTraining2021] in Figure 1. We simulate the pre-training process on two scenarios: (a) standard pre-training on the balanced ImageNet dataset [russakovskyImagenetLargeScale2015]; and (b) drift pre-training on ImageNet-LT [liuLargeScaleLongTailedRecognition2019], which simulates a non-stationary stream with an increasingly imbalanced distribution. For a fair comparison, the feature embeddings are projected using t-SNE and evaluated on a balanced test set. Category colors range from dark to light, representing the transition from tail classes (few training samples) to head classes (abundant samples). As shown in Figure 1, the feature space under long-tailed drift becomes highly distorted: tail categories collapse into a compact region, while head categories dominate most of the embedding manifold. This imbalance contrasts sharply with the well-structured and evenly distributed feature space obtained under balanced pre-training. Such degradation can be attributed to the bias accumulation in the momentum-updated teacher network, which is repeatedly exposed to head-class samples. As the distribution drifts, the teacher provides outdated or skewed supervision, leading to suboptimal representation learning for the under-represented classes.
Building the above analysis, we formalize the interplay among these components using a Structural Causal Model (SCM), illustrated in Figure 2(a). It elucidates the causal relationships among input samples $X$, predictions $Y$, the latent concept drift $D$ inherent in the data stream, and the bias $B$ accumulated in the momentum-updated teacher network [pearl2014interpretation]. Our causal analysis (detailed in Section 2.1) reveals that concept drift $D$ acts as a critical confounder. It not only influences the input features $X$ ($D\rightarrow X$) but also shapes the bias $B$ within the teacher network $(X,D)\rightarrow B$, which in turn affects $Y$, i.e., $(X,B)\rightarrow Y$. This induces a spurious backdoor path $X\leftarrow D\rightarrow B\rightarrow Y$ that can mislead the learning process by inducing false correlations between $X$ and $Y$. Furthermore, the mediation [pearl2022direct] path $X\rightarrow B\rightarrow Y$ confounds the direct contribution of $X\rightarrow Y$, as detailed in Section 2.1.
These causal insights underscore the limitations of standard contrastive methods in drifting environments. Therefore, to counteract the detrimental effects of confounding and biased mediation, we propose a novel pre-training method, Resilient Contrastive Pre-training (RCP). RCP leverages causal intervention techniques [pearl2016causal] to actively mitigate the bias $B$ propagated by concept drift $D$, thereby enabling more robust and generalizable representation learning from non-stationary data streams. In summary, our paper mainly makes the following contributions:
@@
    We are pioneers in uncovering how concept drift impacts contrastive pre-training, specifically identifying the accumulation of bias through the momentum update mechanism. This foundational understanding paves the way for future research into robust pre-training on large-scale, dynamic datasets.
@@
    Leveraging insights from our structural causal model, we propose resilient contrastive pre-training (RCP), a novel method that employs causal intervention to mitigate drift-induced bias. RCP is designed to be straightforward to implement and scalable, making it suitable for pre-training large models on evolving data streams.
@@
    Extensive experiments demonstrate the superiority of RCP across diverse downstream tasks, e.g., long-tailed classification, OOD detection, and domain adaptation, using both fine-tuning and linear probing evaluation protocols. Notably, our analysis of inter-category distances in the learned feature space confirms that RCP effectively mitigates the detrimental effects of drift, leading to more discriminative and robust representations for large-scale pre-training.
@@ 2 Methodology
2.1 Causal Lens on Contrastive Drift Pre-training
Contrastive Drift Pre-training (CDP). Concept drift is a statistical phenomenon wherein the joint distribution $P(X,Y)$ of features $X$ and target variables $Y$ evolves over time [luLearningConceptDrift2019]. We define Drift Pre-training as a pre-training paradigm operating on a data stream $\mathcal{S}_{0,T}=\{S_{0},S_{1},\dots,S_{T}\}$. Each element $S_{t}=\{(x_{i},y_{i})\}_{i=1}^{n_{t}}$ consists of $n_{t}$} samples with feature vector $x_{i}$ and an associated contrastive label $y_{i}$ at timestamp $t$. This timestamp $t$ corresponds to an iteration in the pre-training process. At any given iteration $t$, data samples $\{X,Y\}\in\{(x_{i},y_{i})\}_{i=1}^{n_{t}}$ are drawn from a joint distribution $P_{t}(X,Y)$. Concept drift within this pre-training context is then formally characterized by:
$\exists t\in[0,T)\text{:}P_{t}(X,Y)\neq P_{t+1}(X,Y),$ (1)
where the joint probability $P_{t}(X,Y)$ can be decomposed as $P_{t}(X,Y)=P_{t}(X)\times P_{t}(Y|X)$. Specifically, tailed drift (or long-tailed imbalance) primarily reflects variations in the marginal distribution $P_{t}(X)$, where the frequency of certain classes or patterns changes over time. In contrast, out-of-distribution (OOD) drift corresponds to alterations in the conditional distribution $P_{t}(Y|X)$, indicating that the mapping from features to labels evolves with new or redefined concepts. Both phenomena are manifestations of concept drift, where either the data distribution or the underlying labeling function (or both) changes unpredictably over time. In this work, we aim to develop a pre-training paradigm that can explicitly account for and adapt to these coupled forms of drift, enabling contrastive models to remain robust and consistent across dynamic environments.
CDP investigates contrastive learning under such non-stationary conditions, typically employing momentum-updated teacher-student frameworks [heMomentumContrastUnsupervised2020], i.e.,
$\theta^{m}_{t}=\lambda\theta^{m}_{t-1}+(1-\lambda)\theta^{g}_{t-1},$ (2)
where $\theta^{m}$ signifies the parameters of the teacher network, $\theta^{g}$ denotes those of the student network, and $\lambda\in[0,1)$ represents the momentum coefficient. Momentum $\lambda\theta^{m}_{t-1}$ induces a slower evolution of the teacher network $\theta^{m}$ relative to the student network $\theta^{s}$. While high momentum (e.g., $\lambda=0.999$ [chenEmpiricalStudyTraining2021]) stabilizes training on stationary data, it poses a significant challenge under concept drift: the slowly updating teacher network $\theta^{m}$ can become severely misaligned with the current data distribution $P_{t}(x,z)$, thereby accumulating stale biases from past distributions and hindering the model’s ability to adapt to new or evolved concepts present in the current iteration $S_{t}$.
(a) SCM
@@ 2.1 Causal Lens on Contrastive Drift Pre-training
Objective. Therefore, CDP aims to enable the learning system to synchronize with the evolving data distribution, ensuring that the pre-training model accurately reflect the current state of the data. We formalize this objective as maximizing the agreement between the student model and momentum model over a prospective adaptation window $t,t+\tau]$:
$\max_{\{g_{j}^{\theta}\}_{j=t}^{t+\tau}}\sum_{j=t}^{t+\tau}\mathcal{L}_{\text{agree}}(g_{j}^{\theta}(\tilde{X}_{j}),m_{j}^{\theta}(\hat{X}_{j}))$ (3)
where $\tilde{X}_{j}$ and $\hat{X}_{j}$ symbolize different augmented samples obtained from ${X}_{j}$, $g_{j}^{\theta}$ denotes the student model trained by the data stream $S_{j-\tau,j-1}$ from the drift adaption window with the size of $\tau$, and $m_{j}^{\theta}$ represents the momentum model. And the pre-training is driven by the target metric $\mathcal{L}_{\text{agree}}$ to continuously adapt the drift in a given drift adaptation window of $t,t+\tau]$ time period. This is a standard objective for continuous contrastive learning but fails to explicitly handle the accumulated bias $B$ caused by drift $D$, which we formalize next.
Causal Analysis of CDP. To further understand the challenges posed by concept drift in CDP, we employ a Structural Causal Model (SCM) [pearl1995causal, pearl2016causal]. As shown in Figure 2, the constructed causal graph of $\{X,Y,D,B\}$ presents the following causal connections:
@@
This structure reveals two detrimental pathways that compromise the true causal link $X\rightarrow Y$:
Backdoor Path (Spurious Correlation): $X\leftarrow D\rightarrow B\rightarrow Y$. Concept drift $D$ acts as a confounder. This path introduces a spurious correlation, meaning $X$ and $Y$ appear related solely due to their mutual dependency on $D$ and $B$. For instance, in long-tailed drift, tail samples might be spuriously associated with head categories due to the overwhelming head-class bias $B$.
Mediation Path (Weakening Effect): $X\rightarrow B\rightarrow Y$. The bias $B$ acts as a mediator. This path conveys the drift effect, effectively capturing and weakening the direct and desirable impact of the true feature $X$ on the prediction $Y$.
(a) Memory Bank
(b) MoCo
(c) Our Resilient Contrastive Learning
Figure 3: Comparison between different contrastive pre-training paradigms. (a) The key representations are sampled from a memory bank. (b) A momentum-updated encoder maintains the queue of keys. (c) The workflow of our resilient contrastive pre-training under concept drift streaming. Within the data streaming, a large batch size is opted for a wider drift adaptation window sliding to adapt changes in data distribution. Undergoes various random augmentations, the transformed instances from the identical sample are feature-extracted by both the encoder and the momentum encoder to get the key and value, respectively. A head is utilized to debias the drift and obtain the query of the encoder features. Subsequently, causal intervention is utilized to alleviate concept drift in the data stream within the adaptation window, resulting in the acquisition of two objects for contrastive learning.
@@ 2.2 Causal Interventional Contrastive Objective
To eliminate the confounding influence of drift $D$ (the backdoor path) and recover the genuine causal effect $X\rightarrow Y$, we employ a causal intervention. Drawing from the aforementioned causal relationships, it can be argued that current contrastive pertaining approaches involving the InfoNCE loss [oordRepresentationLearningContrastive2019] essentially leverages the likelihood $P(Y|X)$ to drive the whole network, which is vulnerable by the concept drift confounder $D$ and leads to the emergence of spurious correlations. By marginalizing over the discrete drift contexts $D=\{d_{1},\dots,d_{|D|}\}$ [tangLongTailedClassificationKeeping2021], $P_{t}(Y|X)$ can be expressed as:
$P_{t}(Y|X)=\sum_{i=1}^{|D|}P_{t}(Y|X,B=h(X,d_{i}))P_{t}(d_{i}|X)$ (4)
where $B=h(X,d)$ represents the sample bias $B$ conditioned on features $X$ and a specific drift context $d_{i}$. Following the [dengComprehensiveKnowledgeDistillation2021], $d_{i}$ is simplified to the drift of the latent category center underlying the pre-training. The term $P_{t}(d|X)$ highlights how $X$ might carry information about $d$.
Inspired by recent advancements in applying causal inference to deep learning [lvCausalityInspiredRepresentation2022, choiC2LCausallyContrastive2022, rohekarCausalInterpretationSelfAttention2023], we employ causal intervention $P_{t}(Y|\text{do}(X))$ to cut off the pathway of $D\rightarrow X$ at timestamp $t$ in drift data stream as illustrated in Fig. 2(b), where $\text{do}(\cdot)$ denotes the interventional operation [pearl2016causal, pearl2000models]. In accordance with the backdoor adjustment, the de-confounding operation $\text{do}(X)$ involves stratifying the confounder into discrete segments $D=\{d_{1},d_{2},\ldots,d_{|D|}\}$, thereby rendering $D$ no longer a confounder between $X$ and $Y$. Thus, $P_{t}(Y|\text{do}(X))$ can be expressed as:
$P_{t}(Y|\text{do}(X))=\sum_{i=1}^{|D|}P_{t}(Y|X,B=h(X,d_{i}))P_{t}(d_{i})$ (5)
In Eq. (5), $P_{t}(Y|\text{do}(X))$ compels $X$ to equally consider each confounding factor $d_{i}$ and integrates them collectively to predict $Y$. Then, the original classifier $P_{t}(Y|X)$ is replaced by $P_{t}(Y|\text{do}(X))$ at the iteration $t$, as shown:
$P_{t}(Y|\text{do}(X))\triangleq P_{t}(Y|X)$ (6)
The de-confound classifier mitigates the confounding influence and captures the genuine causality from $X$ to $Y$, thereby enhancing the quality of the contrastive pre-training. Nevertheless, Eq. (5) necessitates costly sampling to approximate $P(Y|\text{do}(X))$ when implementing it in contrastive pre-training, resulting in impractical training times. Fortunately, inspired by prior studies that approximate Eq. (5) in a feed-forward manner, we follow this line of work and employ a single-pass estimation to approximate the intervention effect:
$P_{t}(Y|\text{do}(X))\approx P_{t}(Y|X,B=\sum_{i=1}^{|D|}h(X,d_{i})P_{t}(d_{i}))$ (7)
$\mathcal{L}_{\text{agree}}(g_{j}^{\theta}(\tilde{X}_{j}),m_{j}^{\theta}(\hat{X}_{j}))=P_{t}(Y|X,\sum_{i=1}^{|D|}h(X,d_{i})P_{t}(d_{i}))$ (8)
$=P_{t}(Y|g(\tilde{X}_{j}),\text{Softmax}(h(g_{j}^{\theta}(\tilde{X}_{j}m_{j}^{\theta}(\hat{X}_{j})))),$
where the output of the final expression serves as the similarity score $\mathcal{L}_{\text{agree}}$ for the contrastive loss.
2.3 Resilient Contrastive Pre-training
This section details the concrete architecture and mechanism used to instantiate the causal objective (Eq. (8)).
Table 1: Evaluation results of fine-tuning on long-tailed classification tasks with ImageNet-LT [liuLargeScaleLongTailedRecognition2019] and iNatualist2018 [vanhornINaturalistSpeciesClassification2018a]. The best-performing contrastive pre-training results are highlighted in red. Many, Medium and Few denote the evaluated splits of many-shot (>>100 training samples), medium-shot (20-100 samples) and few-shot (<<20 samples). Top-1 accuracy is applied to evaluate the performance of different methods. $\dagger$ denotes methods that are adjusted for the long-tailed fine-tuning.
ImageNet-LT iNaturalist2018
Methods Venue Backbones Many Medium Few All Many Medium Few All
Training from scratch
cRT [kangDecouplingRepresentationClassifier2019] ICLR’19 61.8 46.2 27.3 49.6 69.0 66.0 63.2 65.2
LWS [kangDecouplingRepresentationClassifier2019] ICLR’19 60.2 47.2 30.3 49.9 65.0 66.3 65.5 65.9
RIDE [wangLongtailedRecognitionRouting2020] ICLR’20 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6
PaCo [cuiParametricContrastiveLearning2021] ICCV’21 ResNet-50 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2
Masked Image Modeling Pre-training
MAE [heMaskedAutoencodersAre2022] CVPR’22 74.7 48.2 19.4 54.5 79.6 70.8 65.0 69.4
LiVT $\dagger$ [xuLearningImbalancedData2023] CVPR’23 ViT-B/16 73.6 56.4 41.0 60.9 78.9 76.5 74.8 76.1
Contrastive Learning Pre-training
BYOL [grillBootstrapYourOwn2020] NeurIPS’20 30.6 8.6 0.9 15.9 28.4 22.6 24.5 24.0
ViT [dosovitskiyImageWorth16x162021] NeurIPS’21 50.5 23.5 6.9 31.6 65.4 55.3 50.9 54.6
DINO [caronEmergingPropertiesSelfSupervised2021] ICCV’21 64.8 35.5 11.3 43.3 76.7 67.1 61.2 65.5
DeiT [touvronDeiTIIIRevenge2022] ECCV’22 70.4 40.9 12.8 48.4 72.9 62.8 55.8 61.0
MoCo v3 [chenEmpiricalStudyTraining2021] ICCV’21 ViT-B/16 70.8 40.7 14.3 48.5 76.6 65.8 62.8 65.6
DINO v2 [zhouDINOWMWorldModels2024] ICML’25 68.4 36.9 11.7 45.4 71.9 62.7 58.2 61.6